Can the following be the lengths of the sides of a right triangle? 15,20,25 Yes or no
Found 4 solutions by ikleyn, josgarithmetic, mananth, Alan3354:
Answer by ikleyn(52800)

.

Apply the Pythagoras, and you will know.



Answer by josgarithmetic(39620)
Answer by mananth(16946)

15, 20, 25
the longest side in a right triangle here is 25
15^2 +20 ^ = 225 +400 = 625
25^2 =625
15^2 +20^2 = 625
By Pythagoras theorem the triangle is a right angled triangle
Yes they are


Answer by Alan3354(69443)

Can the following be the lengths of the sides of a right triangle? 15,20,25 Yes or no
----------------
It's the famous 3-4-5 triangle multiplied by 5.